# A Proactive Decoy Selection Scheme for Cyber Deception using MITRE ATT&CK

### v1

@@ 6.2 Decoy selection performance analysis
We evaluate the various schemes according to the aforementioned metrics. In particular, unless differently stated, the optimal scheme is configured with $\beta=1$ (i.e., mitigation-unaware), the group scheme is computed with $\gamma=0$ and $\rho=1$ (i.e., one group whose techniques can reach all the selected targets).
@@
In Fig. 5 we show the attack path detection ratio achieved by the various schemes according to a variable number of attack targets. The optimal scheme and the predecessor scheme achieve a 100% interception rate of the attack paths for any number of attack targets. This is an expected behavior, as these methods are designed to compute a subset of nodes that disconnect any source in $S$ to any target in $T$. Therefore, the associated selected decoys intercept all attack paths. The performance gap between the optimal and group scheme is explained by the fact that the latter considers only techniques exclusively used by one threat actor and it ignores all possible attack paths enabled by other techniques composing the considered threat profile. Conversely, the optimal scheme is designed to comprehensively accounts for any feasible attack paths that allow to reach the various targets. The random scheme provides the worst performance due to the fact that it does not exploit any information provided by the attack graph, highlighting that the performance gain of optimal decoy selection leverages the preconditions shared among the employed techniques to efficiently deny multiple attack paths. Note that increasing performance gain of the random scheme is due to a higher likelihood to randomly intercept attack paths, which indeed increase with the number of targets.
@@
In Fig. 5 we present the number of decoy selected by each scheme. As a general trend, the overall number of decoys increases for all schemes as a higher number of attack targets generates a higher number of attack paths that need to be intercepted. The optimal scheme always selects the lowest number of decoys, thus minimizing the impact on the attack surface of the system due to the deployed decoys. The random scheme also achieves the same performance since it is configured by design to select the same number of decoys as the optimal method. Conversely, the predecessor scheme provides a rapidly-increasing number of decoys as it restricted to select those techniques causing the outcomes chosen as targets. Therefore, the existence of multiple attack targets reachable by many techniques require a large amount of decoys. This behavior makes such scheme unpractical when the number of considered targets is high due to its poor scalability performance. Finally, the group scheme produces a considerable number of decoys even for a lower number of targets but it less sensitive when this value increases. This depends on the flexibility of the techniques used by the associated threat actors as they are likely to be employed for a plethora of different attack scenarios, allowing their reusability for different attack targets. By combining the number of decoys with their attack path interception capability, we highlight the advantage provided by the optimal scheme with respect to the other approaches as it guarantee the highest interception rate using the lowest amount of decoys.
@@
In Fig. 5 we show the unmitigated technique ratio. The optimal method performance is dictated by the $\beta$ parameter, which associates higher weights to the nodes in $\bar{G}$ corresponding to techniques with mitigations. In particular, compared to the setting $\beta=1$ that provides a mitigation-unaware decoy selection, $\beta=2$ successfully selects a higher percentage of techniques without countermeasures to deny the various attack paths. Nonetheless, regardless of the value of $\beta$, the optimal approach outperforms all other schemes by a significant margin, suggesting that the techniques allowing for a more efficient partition of $\bar{G}$ are indeed without mitigation and therefore the associated attack paths are more effectively detected via a proactive defense approach, such as cyber deception. This observation is further supported by the performance of the group and predecessor schemes which are instead biased to select a higher ratio of techniques with mitigation as decoys. Specifically, the group scheme selects techniques forming attack paths successfully executed in real-world systems, whereas the predecessor scheme selects techniques that have an immediate impact (i.e., the technique outcome) on the environment. In both scenarios, these techniques becomes ideal candidates for the design of possible defense countereasures to mitigate them. As a result, the associated decoys are more likely to overlap with existing security mechanisms. Finally, the random scheme achieves a constant performance regardless of the number of targets, as it is agnostic to the available mitigations.
@@ 6.3 Optimal Solution Analysis
Figure 7: Number of type-AND nodes intercepted by the decoys.
@@
In Fig. 7, we show the average number of nodes with type AND preconditions detected by the decoys. This metric quantifies the capability of the decoys computed with the optimal and group schemes to exploit AND preconditions in $\bar{G}$ in order to detect multiple attack paths. The optimal scheme achieves the best performance since it actively takes advantage of the AND logic that is embedded within the problem formulation. As a matter of fact, an attacker needs to satisfy every precondition of an AND node in order to employ the related technique or to achieve the resulting outcome. This dynamic allows reducing the number of required deceptive artifacts since any technique or outcome of type AND can be efficiently disabled by exposing one of its predecessor techniques with a decoy. Conversely, the group scheme is insensible to the node type as it is shown by the evident performance gap.
@@ 00 $1$ Single threat actor reaching all attack targets
In Fig. 9, we show the attack path interception ratio. In general, the group scheme configurations selecting all threat actors (i.e., $\gamma=1$) detects a higher number of attack paths. These configurations select a higher number of decoys compared to the single threat actors approach (i.e., $\gamma=0$) as they select a higher amount of techniques compatible with the attack targets. Similarly, the compatibility between threat actors and attack targets also impacts the attack path interception performance of the group scheme. The selection of threat actors reaching every attack target (i.e., $\rho=1$) tends to ensure a higher ratio of intercepted attack paths since the selected decoys are more likely to detect techniques composing the considered threat profile. However, when the number of target increases, the selection of all threat actors reaching at least one attack target (i.e., $\gamma=1,\rho=0$) ensures a higher ratio of intercepted attack paths compared to selecting threat actors reaching all attack targets (i.e., $\gamma=1,\rho=1$). This counterintuitive behavior is due to the limited number of groups satisfying the target reachability condition. As the number of attack targets increases, there are fewer threat actors that can simultaneously reach each target. This filtering reduces the number of selected decoys since the number of technique to expose is progressively lower. Consequently, the selected decoys are strictly tailored to detect attack paths used by few threat actors only. In other words, this configuration approximates the single-threat actor approach for a high number of targets as the pool of eligible threat actors is composed by few units (i.e., similarly to the configuration $\gamma=1$), thus explaining the comparable performance. Conversely, when the reachability requirement is $\rho=1$, we observe a more stable performance due to the fact that the scheme has a larger pool of threat actors to choose from, which heavily inflates the attack path detection performance due to the large number of selected decoys. However, when combined with the single threat actor configuration, the resulting decoy selection yields the worst results since the limited number of exposed techniques are rarely used by the various attack paths to reach the considered targets.
@@ 7 Conclusion
We proposed a decoy selection scheme to proactively select deceptive artifacts using an adversarial model built from empirical observations of TTPs. Specifically, we first modeled the attacker capabilities by combining the information provided by enterpriseLang, a threat language modelling for enterprise environments, and MITRE ATT&CK framework. The resulting threat model allows to generate attack graphs expressing the various attack paths between a source node, representing the initial attacker foothold, and a set of attack targets, representing the final attacker goals, using real-world attackers techniques. Each attack path is expressed as a subset of ATT&CK techniques that can be exposed by suitable decoys designed to lure attackers employing those strategies. Then, based on this assumption, we formulated an integer linear optimization problem that computes the minimum number of decoys exposing ATT&CK techniques required to reach a set of predefined attack targets. The proposed formulation accounts for the execution precondition of the various techniques to properly optimize the amount of deceptive artifacts. As a result, the selected decoys can effectively intercept attackers before they accomplish those goals. We assessed the performance of our approach by designing different schemes that partially exploit the cyber threat information provided by our model to compute the decoy selection. The proposed optimal scheme outperformed all the considered benchmarks in all metrics. In particular, by taking advantage of logical dependencies between the various techniques, it selects a considerably lower amount of decoy to intercept any attack path directed to a specific set of attack targets.

### v2

@@ 6.2 Decoy selection performance analysis
We evaluate the various schemes according to the aforementioned metrics. In particular, unless otherwise stated, the optimal scheme is configured with $\beta=1$ (i.e., mitigation-unaware), the group scheme is computed with $\gamma=0$ and $\rho=1$ (i.e., one group whose techniques can reach all selected targets).
@@
In Fig. 6, we show the attack path detection ratio achieved by the various schemes according to a variable number of attack targets. The optimal scheme and the predecessor scheme achieve a 100% interception rate of the attack paths for any number of attack targets. This is an expected behavior, as these methods are designed to compute a subset of nodes that disconnect any source in $S$ to any target in $T$. Therefore, the associated selected decoys intercept all attack paths. The performance gap between the optimal and group scheme is explained by the fact that the latter considers only techniques exclusively used by one threat actor and it ignores all possible attack paths enabled by other techniques composing the considered threat profile. Conversely, the optimal scheme is designed to comprehensively accounts for any feasible attack path that allows to reach the various targets. The random scheme provides the worst performance due to the fact that it does not exploit any information provided by the attack graph, highlighting that the performance gain of optimal decoy selection leverages the preconditions shared among the employed techniques to efficiently deny multiple attack paths. Note that increasing performance gain of the random scheme is due to a higher likelihood to randomly intercept attack paths, which indeed increase with the number of targets.
@@
In Fig. 6, we present the number of decoy selected by each scheme. As a general trend, the overall number of decoys increases for all schemes as a higher number of attack targets generates a higher number of attack paths that need to be intercepted. The optimal scheme always selects the lowest number of decoys, thus minimizing the impact on the attack surface of the system due to the deployed decoys. The random scheme also achieves the same performance since it is designed to select the same number of decoys as the optimal method. Conversely, the predecessor scheme provides a rapidly-increasing number of decoys as it restricted to select those techniques causing the outcomes chosen as targets. Therefore, the existence of multiple attack targets reachable by many techniques require a large amount of decoys. This behavior makes such scheme unpractical when the number of considered targets is high due to its poor scalability performance. Finally, the group scheme produces a considerable number of decoys even for a low number of targets but it less sensitive when this value increases. This depends on the flexibility of the techniques used by the associated threat actors as they are likely to be employed for a plethora of different attack scenarios, allowing their reusability for different attack targets. By combining the number of decoys with their attack path interception capability, we highlight the advantage provided by the optimal scheme with respect to the other approaches as it guarantees the highest interception rate using the lowest amount of decoys.
@@
In Fig. 6, we show the unmitigated technique ratio. The performance of the optimal method is dictated by the $\beta$ parameter, which associates higher weights to the nodes in $\bar{G}$ corresponding to techniques with mitigations. In particular, compared to the setting $\beta=1$ that provides a mitigation-unaware decoy selection, $\beta=2$ successfully selects a higher percentage of techniques without countermeasures to deny the various attack paths. Nonetheless, regardless of the value of $\beta$, the optimal approach outperforms all other schemes by a significant margin, suggesting that the techniques allowing for a more efficient partition of $\bar{G}$ are indeed without mitigation and therefore the associated attack paths are more effectively detected via a proactive defense approach, such as cyber deception. This observation is further supported by the performance of the group and predecessor schemes which are instead biased to select a higher ratio of techniques with mitigation as decoys. Specifically, the group scheme selects techniques forming attack paths successfully executed in real-world systems, whereas the predecessor scheme selects techniques that have an immediate impact (i.e., the technique outcome) on the environment. In both scenarios, these techniques become ideal candidates for the design of possible defense countermeasures to mitigate them. As a result, the associated decoys are more likely to overlap with existing security mechanisms. Finally, the random scheme achieves constant performance regardless of the number of targets, as it is agnostic to the available mitigations.
@@ 6.3 Optimal Solution Analysis
Figure 8: Number of AND nodes intercepted by the decoys.
@@
In Fig. 8, we show the average number of AND nodes detected by the decoys. This metric quantifies the capability of the decoys computed with the optimal and group schemes to exploit AND preconditions in $\bar{G}$ in order to detect multiple attack paths. The optimal scheme achieves the best performance since it actively takes advantage of the AND logic that is embedded within the problem formulation. As a matter of fact, an attacker needs to satisfy every precondition of an AND node in order to employ the related technique or to achieve the resulting outcome. This dynamic allows reducing the number of required deceptive artifacts since any technique or outcome of type AND can be efficiently disabled by exposing one of its predecessor techniques with a decoy. Conversely, the group scheme is insensible to the node type as it is shown by the evident performance gap.
@@ 00 $1$ Single threat actor reaching all attack targets
In Fig. 10, we show the attack path interception ratio. In general, the group scheme configurations selecting all threat actors (i.e., $\gamma=1$) detects a higher number of attack paths. These configurations select a higher number of decoys compared to the single threat actor approach (i.e., $\gamma=0$) as they select a higher amount of techniques compatible with the attack targets. Similarly, the compatibility between threat actors and attack targets also impacts the attack path interception performance of the group scheme. The selection of threat actors reaching every attack target (i.e., $\rho=1$) tends to ensure a higher ratio of intercepted attack paths since the selected decoys are more likely to detect techniques composing the considered threat profile. However, when the number of target increases, the selection of all threat actors reaching at least one attack target (i.e., $\gamma=1,\rho=0$) ensures a higher ratio of intercepted attack paths compared to selecting threat actors reaching all attack targets (i.e., $\gamma=1,\rho=1$). This counterintuitive behavior is due to the limited number of groups satisfying the target reachability condition. As the number of attack targets increases, there are fewer threat actors that can simultaneously reach each target. This filtering reduces the number of selected decoys since the number of technique to expose is progressively lower. Consequently, the selected decoys are strictly tailored to detect attack paths used by few threat actors only. In other words, this configuration approximates the single-threat actor approach for a high number of targets as the pool of eligible threat actors is composed by few units (i.e., similarly to the configuration $\gamma=1$), thus explaining the comparable performance. Conversely, when the reachability requirement is $\rho=1$, we observe a more stable performance due to the fact that the scheme has a larger pool of threat actors to choose from, which heavily inflates the attack path detection performance due to the large number of selected decoys. However, when combined with the single threat actor configuration, the resulting decoy selection yields the worst results since the limited number of exposed techniques are rarely used by the various attack paths to reach the considered targets.
@@ 7 Conclusion
We proposed a decoy selection scheme to proactively select deceptive artifacts using an adversarial model built from empirical observations of TTPs. Specifically, we first modeled the attacker capabilities by combining the information provided by enterpriseLang, a threat language modelling for enterprise environments, and MITRE ATT&CK framework. The resulting threat model allows to generate attack graphs expressing the various attack paths between a source node, representing the initial attacker foothold, and a set of attack targets, representing the final attacker goals, using real-world attackers techniques. Each attack path is expressed as a subset of ATT&CK techniques that can be exposed by suitable decoys designed to lure attackers employing those strategies. Then, based on this assumption, we formulated an integer linear optimization problem that computes the minimum number of decoys exposing ATT&CK techniques required to reach a set of predefined attack targets. The proposed formulation accounts for the execution precondition of the various techniques to properly optimize the amount of deceptive artifacts. As a result, the selected decoys can effectively intercept attackers before they accomplish those goals. We assessed the performance of our approach by designing different schemes that partially exploit the cyber threat information provided by our model to compute the decoy selection. The proposed optimal scheme outperformed all the considered benchmarks in all metrics. In particular, by taking advantage of logical dependencies between the various techniques, it selects a considerably lower amount of decoys to intercept all attack paths directed to a specific set of attack targets.